# Eye-tracked Virtual Reality: A Comprehensive Survey on Methods and Privacy Challenges

### v2

@@ V-B Privacy-preserving Eye Tracking for VR
As most of the works that provide formal privacy guarantees add considerable noise to the data, this negatively affects the performance of the utility tasks. When real-time interaction is not needed, and the utility is limited to data mining, privacy protection from formal methods by adding noise is reasonable; however, especially in applications that require a real-time working capability, the amount of noise necessary for a privacy guarantee can significantly deteriorate the user experience.
@@
Considering this, researchers proposed solutions by underlining the importance of real-time and practical use by trying to limit the amount of noise introduced to real-time gaze data while still reducing the risk of re-identification. David-John et al. [348] presented a privacy protection method by utilizing spatial and temporal downsampling of gaze positions along with additive Gaussian noise when streaming the data to decrease the granularity of the eye-tracking data. The authors’ findings show that re-identification rates drop significantly when proposed methods are applied, without a formal guarantee. At the same time, the performance of gaze prediction as a utility task is minimally affected. In another work, Fuhl et al. [349] proposed training reinforcement learning agents by maximizing the rewards for utility tasks (e.g., expertise prediction, document-type classification) and by minimizing them for privacy tasks (e.g., gender detection, person identification) to balance the privacy and utility. The authors’ approach outperforms privacy protection by using GANs and differentially private manipulation; however, their approach works probabilistically, meaning that it does not formally guarantee privacy and is most appropriate when specific risks and adversary models are known and not expected to change.
In addition to the formal and probabilistic approaches, researchers also studied function-specific systems focused on gaze estimation [350, 351]. Bozkir et al. [350] proposed a cryptography-based method utilizing a randomized encoding-based framework in a three-party setup, where one party is identified as a server (e.g., a cloud instance) that trains machine learning models using sensitive eye movement data, and the other two parties provide the sensitive data in a masked way. While raw and sensitive data samples from input parties are not visible to any party except the data owner, the inference time allows a real-time interaction as long as efficient communication between the parties is available. While this work suits real-time interaction for any VR application, the utilized privacy framework is limited to two data-provider parties. In addition, the balance and focus of the privacy and utility trade-off allegedly concern the privacy of the raw eye image data and the computational efficiency of multi-party estimation of gaze directions. If the estimation task is computationally complex or there is no efficient communication between the parties, the proposed privacy-preserving framework may not work in real time, which can deteriorate the user experience significantly.
@@
More recently, Elfares et al. [351] proposed a federated learning approach for appearance-based gaze estimation in the wild using pseudo-gradient optimization. In federated learning [352], the machine learning models are trained in a decentralized manner, implying that sensitive data samples are not distributed but kept locally, preserving data privacy. The authors showed that both in the person-independent setup and in most person-specific setups, their adaptive federated learning approach outperformed vanilla federated averaging [352] for the gaze estimation task. In follow-up work, Elfares et al. [353] combined federated learning and secure multi-party computation for appearance-based gaze estimation, showing that enhanced privacy does not lead to a decrease in gaze estimation accuracies. While neither of the aforementioned works used data collected from VR to train and evaluate their models, Bozkir et al. [350] used synthetic images that are from the UnityEyes framework [94] that are very comparable to eye images obtained from VR HMDs. In contrast, Elfares et al. [351] used a real-world dataset (i.e., MPIIGaze dataset [121]) to estimate the gaze. However, since VR HMDs can already be considered personal devices, federated learning frameworks and processing data in a decentralized way fit well, and the authors’ approach is also directly applicable in VR. The decentralization concept was recently also proposed in an anonymous eye-tracking data collection protocol for VR by eliminating the third parties for data processing and manipulation purposes [354]. While the authors eliminated third-party platforms or entities by using blockchains and smart contracts in their protocol, if the cryptocurrency wallets used to validate data authenticity are somehow linked to a wallet associated with know-your-customer validation, an adversary could potentially identify the user to a certain extent.
@@
Apart from the research on protecting the privacy of eye movements and gaze, a considerable amount of work focused on degrading iris authentication [355, 356], obfuscation [357], and protecting the personal identities [358] when iris textures are involved. To this end, John et al. [355] proposed an approach by utilizing an optical defocus in an eye-tracker setup. The authors found that such defocus causes errors in the range of calibration errors typical of eye trackers. More recently, John et al. [359] analyzed the security-utility trade-off for iris authentication using the optical defocus in the chin-rest setup and found similar results for degradation. The authors indicated that evaluating a similar setup in an immersive VR environment was in the scope of future work. These works primarily aimed to find a balance between the amount of iris authentication degradation and gaze tracking accuracy in the context of privacy-utility trade-off.
@@
In another work, John et al. [356] proposed adding pixel noise to break the iris signature to protect users from spoofing attacks. The authors argued that replacing up to $50\%$ of the pixels in the eye image while keeping the gaze estimation error less than $2.5^{\circ}$ is possible. In further studies, Eskildsen and Hansen found that an optimal method to remove the iris signature without impacting gaze estimation combines an edge-preserving filter with additive noise [357]. While these works did not directly use eye-tracking data collected from VR setups, as VR with HMDs provides a more controlled environment for eye-tracking data collection, we foresee similar results from VR setups as long as the underlying gaze estimation approaches are the same. In the context of VR eye tracking, Chaudhary and Pelz [358] proposed replacing the iris texture regions with synthetic iris templates using a Rubber Sheet Model on the OpenEDS dataset [45], collected using VR HMDs. The authors found that such video manipulations do not degrade the semantic segmentation and pupil detection performance, similar to the findings of John et al. [356] when the privacy-utility trade-off is considered.
@@
In summary, privacy-preserving eye-tracking methods can be clustered into two groups, one concerning the protection of sensitive gaze and eye movement information over time, including the features that are aggregated from the gaze movements, such as fixation durations or saccade rates, and the other concerning the iris obfuscation and degradation. Ultimately, the recommended method depends on the target application and whether the system is running in real-time or being applied offline to a dataset. We note that real-time methods should be tuned to the privacy and utility context of a given application, and it is often challenging to find a good trade-off if formal guarantees are necessary. We also recommend that usability and user experience be explicitly evaluated when new privacy approaches are proposed for eye-tracking data in VR, as they are essential and cannot be sacrificed to achieve high levels of privacy or security.
@@ V-C Summary
Researchers and practitioners utilize eye movements in VR not only for interaction or human behavior understanding but also for person identification and authentication. The unique viewing behaviors of individuals in previous works show that eye movements can be used for authentication purposes to a certain extent, as identification accuracies of such mechanisms are not as high as those of iris-based mechanisms. Despite that, such mechanisms are helpful for 2FA, particularly for in-app purchases or similar tasks in VR. The number of users in respective authentication databases and eye-tracker sampling frequencies are two of the most important factors for an accurate eye movement-based authentication.
Despite the challenges in performance, eye movements can be used for authentication purposes. In addition, as mentioned, different personal characteristics and demographics can also be identified using this data. Therefore, when authentication or identifying such characteristics is not the main purpose, one should consider utilizing privacy-aware methods for eye tracking. Most prior works suggest applying differential privacy, utilizing cryptographic approaches, and decreasing the granularity of the collected eye movement data to preserve the individuals’ and data privacy in VR.
@@ VI Discussion and Future Directions
Taking the wide range of possibilities with eye-tracking data in VR, state-of-the-art methods to preserve the privacy of such data, and suggested best practices by the VR and privacy communities into account, detailed considerations of privacy are needed for eye tracking [336, 4, 23]. We identify three main directions for the eye-tracking community, particularly for privacy, and discuss these in the following subsections.
@@ VI-A Privacy in Social VR and Metaverse
Eye tracking is heavily integrated into discussions regarding privacy in social VR due to the important role of eye movements as a non-verbal cue during social interaction. Social interactions present special considerations for privacy based on context, as expectations vary between interactions in public spaces and interactions with friends or family in private spaces [360]. Potential solutions to establish privacy in social interactions include modulating the degree of accuracy or resolution of eye movements being mapped to a virtual avatar. On the other hand, researchers have found that non-verbal communication, through eye movements and gestures, can help preserve privacy for users who wish to avoid vocal communication in social VR [361], suggesting that users desire control over VR sensor data streams to preserve privacy. Analogous methods exist for modulating privacy in telecommunications today, as users commonly turn off their webcam feed during Zoom or video calls to preserve the privacy of not paying attention or of their background environment. Generating synthetic face and eye animations directly from audio has been proposed [362], which can help preserve privacy from raw eye movements, but is still prone to animation artifacts or incorrectly relaying social cues of the user. Open areas where advancements in understanding privacy for social VR would be critical are environments geared towards children [363], individuals with behavioral conditions such as ASD [364], and communities with privacy norms that vary from the general population [365].
@@
Recent usage of the term “Metaverse” involves a broad vision of the future of immersive reality. While exact definitions vary, the broad futuristic visions include a connected and integrated usage of VR that spans many aspects of our daily lives, superseding the use of the Internet globally today. The privacy concerns introduced by such a deeply connected and immersive VR are yet to be realized, but the long-term implications of capturing eye-tracking data of users in such environments are a critical line of future research. While unexplored in current VR platforms that include functionalities akin to a ‘Metaverse,’ an example of potential privacy concerns for the future is captured by Tadayoshi Kohno’s short story “The Schuhmacher” [366]. The story represents a fictional reality where persistent tracking of behavior for advertising to customers is a two-way street that both benefits a shopkeeper’s business and has negative impacts on his relationship and social status within his town. Privacy concerns due to eye tracking on a societal scale are largely unexplored. To further establish the privacy risks that arise from eye tracking in naturalistic real-world scenarios, eye-tracking datasets such as the Visual Experience Database [367] can be explored. Such databases capture large-scale gaze data from typical daily activities, as opposed to prescribed experimental tasks that comprise the vast majority of research datasets. Characterizing how frequently privacy-sensitive scenarios arise due to eye tracking, and the magnitude of the privacy risk from the users’ perspective, can lead to an understanding of how to anticipate long-term risks introduced by an eye-tracked “Metaverse.”
@@ VI-B Privacy vs Utility and Usability
Computational methods that attempt to preserve the privacy of the users mostly achieve privacy by adding a certain amount of noise, as was the case for the works that utilize differential privacy [25, 356, 312]. However, as reported by these previous works, the utility-privacy trade-off should be taken care of. In differential privacy, it is especially challenging to find a good spot in terms of utility-privacy trade-off considering the eye-tracking datasets, because the longer the signals are, the more noise should be added due to higher sensitivities that contribute to the noise. As eye-tracking data is already a noisy source of information, especially when utilized in real time, adding additional noise to preserve privacy may ruin the user experience. Li et al. [346] addressed this issue by obtaining information on how much users enjoyed the experience when they provided different utility-privacy levels during game play in a desktop setup. With the immersion provided by the VR setups, the user experience and enjoyment levels already change without any privacy provided in the first place, and it is an open question of how the privacy-utility trade-off transfers to immersive setups. In addition, even if the privacy-preserving solutions are used offline, such as private data mining, as high amounts of noise may destroy the patterns in the data, the eye-tracking community needs to find private data representations for different utility tasks (e.g., privacy-preserving scanpath comparison [368]) that have little effect on the utility tasks while preserving privacy.
@@
Privacy-preserving computations aside, providing such solutions in an adaptive and user-centric way is equally important. For instance, one user might be purely interested in a better user experience and utility, while the other one might prefer a privacy-enhanced experience during the use of VR environments. In addition, the particular use case context can also impact the privacy-utility trade-off for a user. For example, in real-time gaze-based interaction for a point-and-shoot game, user subjective experience and performance are critical to consider, in contrast to machine learning-based back-end use case contexts such as gaze-based activity recognition [369]. To achieve this, user privacy behaviors should be known beforehand, and virtual environments can be designed and adapted accordingly. Such privacy aspects, including preferences and attitudes along with usability, have been well studied for a wide range of devices, including smartphones [370], IoT devices [371], lifelogging cameras [372], smart glasses [373], or augmented reality glasses [374]. Despite this, these devices utilize real environments with a negligible to zero amount of vision augmentation. We argue that studying similar user behaviors for VR and HMDs with a focus on eye tracking will enable the deployment of user-adaptive privacy schemes for virtual spaces. Such adaptations should also go hand in hand with privacy regulations since regulations differ based on the countries or regions [375].
@@ VII Conclusion
This paper systematically covered eye-tracking research in VR and the implications of security and privacy, including authentication schemes and privacy-aware methods. To this end, we scanned the scientific literature between 2012 and 2022 in 35 major venues for VR, eye tracking, and privacy. In addition to this literature review, we further provided and discussed three main research directions, especially by keeping privacy as the primary focus, and indicated the importance of privacy, as well as utility and usability.
VIII Acknowledgments
Parts of this work were conducted when EB and EK were with the University of Tübingen. EB and EK acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC number 2064/1 - Project number 390727645. We also acknowledge funding from the DFG - Project number 491966293. The authors used Grammarly to fix grammatical and typographical issues.

### v3

@@ V-B Privacy-preserving Eye Tracking for VR
As most works that provide formal privacy guarantees add considerable noise to the data, this negatively affects the performance of utility tasks. When real-time interaction is not required, and the utility is often limited to data mining, privacy protection through formal methods by adding noise is reasonable. However, especially in real-time applications, the amount of noise necessary for a privacy guarantee can significantly deteriorate the user experience.
@@
Considering this issue, researchers proposed solutions by emphasizing the importance of real-time and practical use, aiming to minimize the amount of noise introduced to real-time gaze data while also reducing the risk of re-identification. [311] presented a privacy protection method by utilizing spatial and temporal downsampling of gaze positions with additive Gaussian noise when streaming the data to decrease the granularity of the eye-tracking data. The findings show that re-identification rates drop significantly when proposed methods are applied, without a formal guarantee, whereas the performance of gaze prediction as a utility task is minimally affected. In another work, [312] proposed training reinforcement learning agents by maximizing rewards for utility tasks (e.g., expertise prediction, document-type classification) and by minimizing them for privacy tasks (e.g., gender detection, person identification) to balance the privacy and utility. This approach outperforms privacy protection by utilizing GANs and differentially private manipulation; however, it also operates probabilistically, meaning that it does not formally guarantee privacy and is most appropriate when specific risks and adversary models are known and not expected to change.
@@
In addition to the formal and probabilistic approaches, researchers also studied function-specific guarantees focused on gaze estimation. [313] proposed a cryptography-based method utilizing a randomized encoding-based framework in a three-party setup, where one party is identified as a server (e.g., a cloud instance) that trains machine learning models using sensitive eye movement data, and the other two parties provide the sensitive data in a masked way, protecting the data privacy. While this work is suitable for real-time interaction in any VR application, the utilized privacy framework is limited to two data-provider parties and requires efficient communication between the involved parties. More recently, [314] proposed a federated learning approach for appearance-based gaze estimation in the wild using pseudo-gradient optimization. In federated learning [315], machine learning models are trained in a decentralized manner, implying that sensitive data samples are not distributed but remain locally, thereby preserving data privacy. The authors demonstrated that their adaptive federated learning approach outperformed vanilla federated averaging [315] in both person-independent setups and most person-specific setups for the gaze estimation task. In follow-up work, [316] combined federated learning and secure multi-party computation for appearance-based gaze estimation, demonstrating that enhanced privacy does not result in a decrease in gaze estimation accuracy. While neither of the works above utilized data collected from VR to train and evaluate their models [313, 314], since VR HMDs can already be considered personal devices, federated learning frameworks and decentralized data processing fit well, and the authors’ approaches are directly applicable in VR. The decentralization concept was recently also proposed in an anonymous eye-tracking data collection protocol for VR by eliminating the third parties for data processing and manipulation purposes [317]. While the authors eliminated third-party entities by using blockchains and smart contracts, wallets associated with know-your-customer validation can still lead to accurate user identification to a certain extent.
@@
Apart from protecting the privacy of eye movements and gaze, a considerable amount of work focused on degrading iris authentication [318, 319], obfuscation [320], and protecting the personal identities [321] when iris textures are involved. To this end, [318] proposed an approach by utilizing an optical defocus in an eye-tracker setup, and found that such defocus causes errors in the range of calibration errors typical of eye trackers. More recently, [322] analyzed the security-utility trade-off for iris authentication using the optical defocus in the chin-rest setup and found similar results for degradation. The authors indicated that evaluating a similar setup in an immersive VR environment was in the scope of future work. These works primarily aimed to find a balance between the degradation of iris authentication and the accuracy of gaze tracking in the context of a privacy-utility trade-off.
@@
In another work, [319] proposed adding pixel noise to break the iris signature to protect users from spoofing attacks, and argued that replacing up to $50\%$ of the pixels in the eye image while keeping the gaze estimation error less than $2.5^{\circ}$ is possible. In further studies, [320] found that an optimal method to remove the iris signature without impacting gaze estimation combines an edge-preserving filter with additive noise. While these works did not directly utilize eye-tracking data collected from VR setups, as VR with HMDs provides a more controlled environment for eye-tracking data collection, we foresee similar results from VR setups, provided the underlying gaze estimation approaches are the same. In the context of VR eye tracking, [321] proposed replacing the iris texture regions with synthetic iris templates using a Rubber Sheet Model on the OpenEDS dataset [37], collected using VR HMDs, and found that such video manipulations do not degrade the semantic segmentation and pupil detection performance, similar to the previous findings [319].
@@
In summary, privacy-preserving eye-tracking methods can be categorized into two groups: one focuses on protecting sensitive gaze and eye movement information over time, including features aggregated from gaze movements, such as fixation durations or saccade rates, and the other focuses on iris obfuscation and degradation. Ultimately, the recommended method depends on the target application and whether the system is running in real-time or being applied offline to a dataset. We note that real-time methods should be tuned to the privacy and utility context of a given application, and it is often challenging to find a good trade-off if formal guarantees are necessary. We also recommend that usability and user experience be explicitly evaluated when new privacy approaches are proposed, as they are essential and cannot be sacrificed to achieve high levels of privacy or security.
@@ V-C Summary
Researchers and practitioners utilize eye movements in VR not only for interaction and understanding human behavior, but also for person identification and authentication. The individuals’ unique viewing behaviors demonstrate that eye movements can be utilized for authentication purposes to a certain extent, as the identification accuracies of such mechanisms are not as high as those of iris-based mechanisms. Despite that, such mechanisms are helpful for 2FA, particularly for in-app purchases or similar tasks in VR. The number of users in respective authentication databases and eye-tracker sampling frequencies are two of the most critical factors for an accurate eye movement-based authentication. Overall, despite the challenges, eye movements can be used for authentication purposes. Additionally, various personal characteristics and demographics can be identified using this data. Therefore, when authentication or identifying such characteristics is not the primary purpose, one should consider utilizing privacy-aware methods for eye-tracking data. Most prior works suggest applying differential privacy, utilizing cryptographic approaches, and decreasing the granularity of the collected eye movement data to preserve the individuals’ and data privacy in VR.
@@ VI Discussion and Future Directions
Taking into account the wide range of possibilities enabled by eye tracking, the state-of-the-art methods for preserving the privacy of such data, and the best practices recommended by the VR and privacy communities, a careful and detailed consideration of privacy is essential for eye tracking [4, 16]. We identify three main directions for the eye-tracking and privacy communities, and discuss these in the following.
@@ VI-A Privacy in Social VR and Metaverse
Eye tracking is heavily integrated into discussions regarding privacy in social VR due to the critical role of eye movements as a non-verbal cue during social interaction. Social interactions present special privacy considerations based on context, as expectations vary between interactions in public spaces and interactions with friends or family in private spaces [323]. Potential solutions to establish privacy in social interactions include modulating the degree of accuracy or resolution of eye movements being mapped to a virtual avatar. On the other hand, researchers have found that non-verbal communication, through eye movements and gestures, can help preserve privacy for users who wish to avoid vocal communication in social VR [324], suggesting that users desire control over VR sensor data streams to protect privacy. Analogous methods exist for modulating privacy in telecommunications today, as users commonly turn off their webcam feed during video calls to preserve the privacy of not paying attention or of their background environment. Generating synthetic face and eye animations directly from audio has been proposed [325], which can help preserve privacy from raw eye movements, but is still prone to animation artifacts or incorrectly relaying users’ social cues. Open areas where advancements in understanding privacy for social VR would be critical include environments geared towards children [326], individuals with behavioral conditions such as ASD [327], and communities with privacy norms that differ from the general population [328].
@@
Recent use of the term “Metaverse” involves a broad vision of the future of immersive reality. While exact definitions vary, the broad futuristic visions include a connected and integrated usage of VR that spans many aspects of our daily lives, superseding the use of the Internet today. The privacy concerns introduced by such deeply connected and immersive VR are yet to be realized, but the long-term implications of capturing eye-tracking data of users in such environments are a critical line of future research. While unexplored in current VR platforms that include functionalities akin to a Metaverse, an example of potential privacy concerns for the future is captured by the short story “The Schuhmacher” [329]. The story represents a fictional reality where persistent behavior tracking for advertising to customers is a two-way street that both benefits a shopkeeper’s business and has negative impacts on his relationship and social status within his town. Privacy concerns due to eye tracking on a societal scale are largely unexplored. To further establish the privacy risks associated with eye tracking in naturalistic, real-world scenarios, eye-tracking datasets such as the Visual Experience Database [330] can be examined. Such databases capture large-scale gaze data from typical daily activities, as opposed to experimental tasks that comprise the vast majority of research datasets. Characterizing how frequently privacy-sensitive scenarios arise due to eye tracking, and the magnitude of the privacy risk from the users’ perspective, can lead to an understanding of how to anticipate long-term risks of the eye-tracked Metaverse.
@@ VI-B Privacy vs Utility and Usability
Computational methods that attempt to preserve the privacy of the users mostly achieve privacy by adding a certain amount of noise, as was the case for the works that utilize differential privacy [18, 319, 279]. However, as reported by these previous works, the privacy-utility trade-off should be taken care of. In differential privacy, it is especially challenging to find a good spot in terms of privacy-utility trade-off, considering the eye-tracking datasets, because the longer the signals are, the more noise should be added due to higher sensitivities that contribute to the noise. As eye-tracking data is already a noisy source of information, especially when utilized in real time, adding additional noise to preserve privacy may ruin the user experience. Li et al. [309] addressed this issue by obtaining information on how much users enjoyed the experience when they provided different privacy-utility levels during game play in a desktop setup. With the immersion provided by the VR setups, the user experience and enjoyment levels already change without any privacy provided in the first place, and it is an open question of how the privacy-utility trade-off transfers to immersive setups. In addition, even if the privacy-preserving solutions are used offline, such as private data mining, as high amounts of noise may destroy the patterns in the data, the eye-tracking community needs to find private data representations for different utility tasks (e.g., privacy-preserving scanpath comparison [331]) that have little effect on the utility tasks while preserving privacy.
@@
Privacy-preserving computations aside, providing such solutions in an adaptive and user-centric way is equally important. For instance, one user might be purely interested in a better user experience and utility, while the other one might prefer a privacy-enhanced experience during the use of VR environments. In addition, the particular use case context can also impact the privacy-utility trade-off for a user. For example, in real-time gaze-based interaction for a point-and-shoot game, user subjective experience and performance are critical to consider, in contrast to machine learning-based back-end use case contexts such as gaze-based activity recognition [332]. To achieve this, user privacy behaviors should be known beforehand, and virtual environments can be designed and adapted accordingly. Such privacy aspects, including preferences and attitudes along with usability, have been studied for other wearable devices, including lifelogging cameras [333], or AR glasses [334]. Despite this, these devices utilize real environments with limited to zero amount of vision augmentation. We argue that studying similar user behaviors for VR and HMDs with a focus on eye tracking will enable the deployment of user-adaptive privacy schemes for virtual spaces. Such adaptations should also align with privacy regulations since regulations differ based on the countries or regions [335].
@@ VII Conclusion
This paper comprehensively covers eye-tracking research in VR and its implications for security and privacy, including authentication schemes and privacy-aware methods. To this end, we scanned the scientific literature from 2012 to 2022 across 35 major venues for VR, eye tracking, and privacy. We further provided and discussed three main research directions, with a primary focus on privacy, and highlighted the importance of privacy, as well as utility and usability.
Acknowledgments
Parts of this work were conducted when EB and EK were with the University of Tübingen, and they acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC number 2064/1 - Project number 390727645 and the DFG - Project number 491966293. The authors used Grammarly to improve readability.